10 In three moles of ethane (C2H6), calculate the following :


(i) Number of moles of carbon atoms.


(ii) Number of moles of hydrogen atoms. 


(iii) Number of molecules of ethane.


Grade:upto college level

3 Answers

Radhika Batra
10 years ago

i) 1 mole of C2H6 contains 2 moles of carbon atoms.

Number of moles of carbon atoms in 3 moles of C2H6

= 2 × 3 = 6

(ii) 1 mole of C2Hcontains 6 moles of hydrogen atoms.

Number of moles of carbon atoms in 3 moles of C2H6

= 3 × 6 = 18

(iii) 1 mole of C2Hcontains 6.023 × 1023 molecules of ethane.

Number of molecules in 3 moles of C2H6

= 3 × 6.023 × 1023 = 18.069 × 1023

prince parihar
10 years ago

1)c2h6 contain 2 carbon atom so in three moles of ethane it will contain 3*2=6 

2)c2h6 contain 6 carbon aton so in three moles of ethane it will contain 6*3=18

3)1 mole contain 6.022*1023 so 3 mole will contain 6.022*1023*3=18.069*1023

Rakesh Singh
10 years ago

Tough question !